OpenStudy (anonymous):

Milan uses a probability simulator to roll a six-sided number cube 100 times and to flip a coin 100 times. The results of the experiment are shown below: Number on the Cube Number of Times Rolled 1 10 2 8 3 33 4 29 5 11 6 9 Heads Tails 29 71 Using Milan's simulation, what is the probability of rolling a 5 on the number cube and the coin landing on tails? fraction 11 over 10,000 fraction 71 over 10,000 fraction 82 over 10,000 fraction 781 over 10,000

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

@Notamathgenius

OpenStudy (anonymous):

@Keigh2015

OpenStudy (anonymous):

plz help. im desperate.

OpenStudy (keigh2015):

I would say the first one I guess

OpenStudy (anonymous):

\[\frac{11}{100}*\frac{71}{100}=\frac{781}{10000} \]

OpenStudy (anonymous):

OMG THAT MAKES SENSE THANK YOU! thanks for not just giving the answer as well hhehehehehe

OpenStudy (anonymous):

Your welcome and thank you for the medal.